# Subgradient Convergence Implies Subdifferential Convergence on Weakly Convex Functions: With Uniform Rates Guarantees

### v1

@@ Abstract
In nonsmooth, nonconvex stochastic optimization, understanding the uniform convergence of subdifferential mappings is crucial for analyzing stationary points of sample average approximations of risk as they approach the population risk. Yet, characterizing this convergence remains a fundamental challenge.
This work introduces a novel perspective by connecting the uniform convergence of subdifferential mappings to that of subgradient mappings as empirical risk converges to the population risk. We prove that, for stochastic weakly-convex objectives, and within any open set, a uniform bound on the convergence of subgradients—chosen arbitrarily from the corresponding subdifferential sets—translates to a uniform bound on the convergence of the subdifferential sets itself, measured by the Hausdorff metric.
Using this technique, we derive uniform convergence rates for subdifferential sets of stochastic convex-composite objectives. Our results do not rely on key distributional assumptions in the literature, which require the population and finite sample subdifferentials to be continuous in the Hausdorff metric, yet still provide tight convergence rates. These guarantees lead to new insights into the nonsmooth landscapes of such objectives within finite samples.
@@ 1 Introduction
At the core of nonsmooth, nonconvex stochastic optimization lies a fundamental yet underexplored challenge—characterizing the statistical uniform convergence behavior of set-valued subdifferential mappings [SX07]. This convergence is essential for understanding how the empirical risk, derived from sample average approximation, approximates the true population risk. While results on the uniform convergence of objective values provide important insights, they do not directly characterize the convergence behaviors of stationary points formulated via subdifferentials. Analyzing the uniform convergence of these set-valued subdifferential mappings of the empirical risk is crucial for a comprehensive understanding of statistical estimation and the behaviors of numerical algorithms in stochastic programming [SDR21, Section 7.2]. Despite its importance, this characterization of subdifferential set convergence remains elusive, warranting further investigation (see Section 1.1 for related works).
@@
This work addresses this gap by introducing a new perspective to understand the uniform convergence behavior of subdifferential sets, specifically for stochastic weakly convex objectives—a prominent class of nonsmooth, nonconvex objectives with widespread applications [PR92, PR96, DD20]. Our approach hinges on a fundamental property of weakly convex functions. Specifically, for any pair of weakly convex functions, we establish that over any arbitrary open set within their domains, the uniform bound on the distance between their respective set-valued subdifferential mappings, measured via the Hausdorff metric, is upper bounded by the uniform bound on the distance between any corresponding vector-valued subgradient mappings of these functions (Theorem 3.1).  This deterministic result immediately implies a crucial connection between the uniform convergence of subdifferentials and the uniform convergence of subgradient mappings in sample average approximations, enabling the analysis of the former through the latter (Theorem 3.4).
@@
We further apply our technique to an important subclass of stochastic weakly-convex objectives called stochastic convex-composite objectives [Bur85, DR18], which include robust phase retrieval [DR19], blind deconvolution [CDDD21], matrix sensing [LZMCSV20, CCD+21], conditional Value-at-Risk [RU00], and nonlinear quantile regressions [KP96], commonly encountered in the field of machine learning, statistics, imaging, and risk management (Theorem 3.9). We establish uniform convergence of subdifferential sets of stochastic weakly-convex objectives by analyzing the convergence behavior of stochastic subgradient mappings (Theorem 4.2). Our result, which does not rely on a common distributional assumption in the literature requiring the subdiffrential mappings to be continuous under Hausdorff metric [SX07, Theorem 3], provides tight convergence rates in the context of many existing studies (see Section 1.1 for a detailed explanation). For instance, in robust phase retrieval [DR19, DDP20], our results (Theorem 5.1) achieve a tight $\sqrt{d/m}$ rate (modulo logarithmic factors on $d,m$) for uniform convergence of subdifferential mappings under the Hausdorff distance where $d$ is the dimension and $m$ is the sample size. This is complementary to the state-of-the-art $4]{d/m}$ rate, which is established under the graphical distance of subdifferentials, through a different technique based on Attouch’s epigraphical convergence theorem [Att77, DD22]. This leads to new understanding of the nonsmooth landscape of these objectives, e.g., the landscape of the finite sample robust phase retrieval initially studied in [DDP20], including the locations and convergence properties of stationary points.
@@ 1.1 Related Work and Contributions
A significant body of literature has utilized set-valued variational analysis [RW98] and random set theory [Mol05] to study the uniform convergence of subdifferentials from empirical to population risks [SX07]. This technique is popular because subdifferential mappings of many nonsmooth and nonconvex functions—including weakly convex functions—often exhibit structural properties such as outer semicontinuity, and the subdifferentials themselves form convex sets, unlike discontinuous subgradient mappings [RW98]. This technique discretizes the parameter space, achieving initial subdifferential convergence at discrete points via Artstein and Vitale’s strong law of large numbers for random convex sets [AV75]. The core assumption is then the subdifferential mapping of both population and the finite sample risks are continuous under the Hausdorff metric. This continuity—not merely outer semicontinuity—is crucial, see [SX07, Remark 2], as it enables the extension of the initial convergence from discrete points to uniform convergence of subdifferential sets over the entire parameter space. Without continuity, using outer-semicontinuity this approach can still ensure uniform convergence of enlarged empirical subdifferentials to enlarged population subdifferentials; however, this does not apply to the original subdifferentials [SX07].
@@
Therefore, a common and core assumption of this approach to attain uniform convergence is the continuity needed for both the population and empirical subdifferentials; while it facilitates many applications, it is sometimes restrictive and does not capture the full picture of subdifferential convergence [SX07, Remark 2]. For achieving uniform convergence with a rate guarantee, stronger assumptions, such as Hölder continuity of subdifferentials or related $H$-calmness conditions, are often necessary [Xu10]. Some research aims to relax the requirement to piecewise continuity of subdifferential sets over disjoint regions, but this still necessitates the underlying probability measure being nonatomic to ensure that samples almost surely avoid the discontinuity boundaries of the subdifferentials [RX11].
@@
This work presents a novel perspective to establish uniform convergence of subdifferential mappings that distinguishes it apart from the existing literature. Specifically, our work targets the domain of stochastic weakly-convex classes, pivotal to numerous data-driven applications [DD20]. The fundamental principle of our approach is that, for stochastic weakly convex objectives, achieving uniform convergence—and the associated rates—of subdifferential sets using sample average approximation requires only the uniform convergence and rates of any chosen pair of the subgradient mappings of the corresponding empirical and population risk functions. As a result, we can derive uniform convergence (and also rates) for a specific subclass termed stochastic convex-composite objectives, without the need for continuity assumptions on the subdifferential mappings of either population or finite sample objectives often required in the existing approach. Importantly, obtaining uniform convergence for discontinuous subgradient mappings in stochastic convex-composite objectives still demands sophisticated tools from statistical learning theory, such as chaining, the concept of shattering, and the notion of Vapnik–Chervonenkis dimension [Vap13].
@@
Notably, there is a recent advance in the literature that also targets at establishment of uniform subdifferential convergence within the domain of stochastic weakly convex classes [DDP20, DD22]. However, there are significant differences between our results and their results. First, there is a subtle difference in terms of the criteria for convergence. Our notion of uniform convergence on any open domain implies their notion of graphical convergence of subdifferential mappings on the same domain, but the reverse is not generally true (see Remark of Theorem 3.1). Therefore, our results are established under a topologically weaker notion of convergence. Second, there is an important difference in the approaches and the established rates. Their approach, grounded in the utilization of Moreau envelope smoothing [Mor65], and Attouch’s epigraphical convergence theorem [Att77], achieves a convergence rate whose dependence on the sample size $m$ is often at $m^{-1/4}$. The reason is that the core principle behind their approach, which aims to derive graphical subdifferential distance bounds through the closeness of objective values, often inherently results in suboptimal guarantees in terms of the sample size [DD22, Section 5]. Our results are notably precise regarding the sample size $m^{-1/2}$, although they apply mainly to stochastic convex-composite objectives and introduce additional logarithmic factors. This indicates that further research is essential to fully delineate the advantages and limitations of both approaches.
@@
Finally, there are many interesting alternative ideas in the literature based on directly smoothing the nonsmooth objective to attain uniform convergence of gradients of the smoothed objectives [XZ09], also used extensively in the statistics literature, e.g., in the context of quantile regressions [Hor98]. These results, however, do not directly characterize the convergence behavior of the original subdifferential sets.
@@ 2 Notation and Basic Definitions
For a function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, its epigraph is the set $\{(x,\alpha)\in\mathbb{R}^{d}\times\mathbb{R}\mid\alpha\geq f(x)\}$. It is called closed if its epigraph is closed in $\mathbb{R}^{d}\times\mathbb{R}$. It is called subdifferentially regular at $x$ if $f(x)$ is finite and its epigraph is Clarke regular at $(x,f(x))$ as a subset of $\mathbb{R}^{d}\times\mathbb{R}$ [RW98, Definition 7.25]. A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is said to be $\mathcal{C}^{1}$ smooth on $\mathbb{R}^{d}$ if it has continuous gradients $\nabla f:\mathbb{R}^{d}\to\mathbb{R}^{d}$.
@@ 2.1 Weakly Convex Functions and Subdifferentials
We say $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is $\lambda$ locally weakly-convex near $x$ (also known as lower-$\mathcal{C}^{2}$ [RW98] or semiconvex [BDLM10]) if there exists $\epsilon>0$ such that
@@
is convex [RW98, Chapter 10.G]. We say a function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is a locally weakly convex function on $\mathcal{O}\subseteq\mathbb{R}^{d}$ if at every $x\in\mathcal{O}$, there is $\lambda<\infty$ such that $f$ is $\lambda$ locally weakly-convex near $x$. For a function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ and a point $x$ with $f(x)$ finite, we let $\partial f(x)$ denote the Fréchet subdifferential (or regular [RW98, Chapter 8.B]) of $f$ at the point $x$,
@@
Here, any $g\in\partial f(x)$ is referred to as a subgradient of $f$ at the point $x$. We say $v$ is a horizon subgradient of $f$ at $x$ with $f(x)$ finite, written as $v\in\partial^{\infty}f(x)$ if there exist sequences $x_{i},v_{i}\in\partial f(x_{i})$ and $\tau_{i}\to 0$ satisfying $(x_{i},f(x_{i}),\tau_{i}v_{i})\to(x,f(x),v)$ [RW98, Chapter 8.B].
Notably, when $f$ is smooth, the subdifferential $\partial f(x)$ consists only of the gradient $\{\nabla f(x)\}$, while for convex function, it coincides with the subdifferential in convex analysis [Roc70]. For a weakly convex function $f$, the subdifferential $\partial f(x)$ is non-empty, compact, and convex for every $x$ in the interior of ${\rm dom}f$. Notably, these properties follow from the corresponding results for convex function in convex analysis [RW98, Chapter 8].
@@ 2.2 Set-valued Analysis
$\liminf_{n}A_{n}=\{y:\exists y_{n}\in A_{n}~{}\text{s.t.}~{}y_{n}\to y\}.$
@@
For every weakly convex function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, the subgradient mapping $\partial f:{\rm int}~{}{\rm dom}(f)\rightrightarrows\mathbb{R}^{d}$ is outersemicontinuous [RW98, Chapter 8].
@@ 2.3 Probability Space and Random Set-Valued Map
Let $(\Xi,\mathscr{G},\mathbb{P})$ denote a probability space. We frequently consider random set-valued mapping of the form $\mathcal{A}:X\times\Xi\rightrightarrows\mathbb{R}^{d}$. Let $\mathscr{B}$ denote the space of nonempty, compact subsets of $\mathbb{R}^{d}$. By [RW98, Theorem 14.4], we say the mapping $\xi\mapsto\mathcal{A}(x,\xi)$ is measurable if and only if for every $\mathcal{S}\in\mathscr{B}$, $\mathcal{A}(x,\cdot)^{-1}\mathcal{S}=\{\omega:A(x,\omega)\in\mathcal{S}\}$ is a measurable set under $(\Xi,\mathscr{G},\mathbb{P})$.
By a selection of the random set $\mathcal{A}(x,\xi)$, we refer to a random vector $a(x,\xi)\in\mathcal{A}(x,\xi)$, meaning that $\xi\mapsto a(x,\xi)$ is measurable. Note that such selection exists if $\xi\mapsto\mathcal{A}(x,\xi)$ is closed-valued and measurable [Aum65], see also [AF09, Theorem 8.4]. For a map $\mathcal{A}:X\times\Xi\rightrightarrows\mathbb{R}^{d}$ and a probability measure $\mathbb{P}$, we define its expectation, following Aumann [Aum65]. At every $x\in X$:
@@
We say $f:\mathbb{R}^{d}\times\Xi\to\mathbb{R}\cup\{+\infty\}$ is a random function when $\xi\mapsto f(x,\xi)$ is measurable at every $x\in X$ [SDR21, Chapter 7]. Suppose a random function $f$ also satisfies $x\mapsto f(x,\xi)$ is locally weakly convex, and real-valued near a point $x\in\mathbb{R}^{d}$, then the subdifferential $\partial f(x,\xi)$ satisfies $\xi\mapsto\partial f(x,\xi)$ is measurable at that $x\in\mathbb{R}^{d}$, following [SDR21, Section 7.2.6] and [XZ09, Proposition 2.1].

### v2

@@ Abstract
….. In nonsmooth, nonconvex stochastic optimization, understanding the uniform convergence of subdifferential mappings is crucial for analyzing stationary points of sample average approximations of risk as they approach the population risk. Yet, characterizing this convergence remains a fundamental challenge. This work introduces a novel perspective by connecting the uniform convergence of subdifferential mappings to that of subgradient mappings as empirical risk converges to the population risk. We prove that, for stochastic weakly-convex objectives, and within any open set, a uniform bound on the convergence of subgradients—chosen arbitrarily from the corresponding subdifferential sets—translates to a uniform bound on the convergence of the subdifferential sets itself, measured by the Hausdorff metric. Using this technique, we derive uniform convergence rates for subdifferential sets of stochastic convex-composite objectives. Our results do not rely on key distributional assumptions in the literature, which require the population and finite sample subdifferentials to be continuous in the Hausdorff metric, yet still provide tight convergence rates. These guarantees lead to new insights into the nonsmooth landscapes of such objectives within finite samples.
keywords:
Subdifferential, Subgradient, Uniform Convergence, Population Risk, Sample Average Approximation, Weak Convexity, Hausdorff Metric.
@@ 1 Introduction
At the core of nonsmooth, nonconvex stochastic optimization lies a fundamental yet underexplored challenge—characterizing the statistical uniform convergence behavior of set-valued subdifferential mappings [1]. This convergence is essential for understanding how the empirical risk, derived from sample average approximation, approximates the true population risk. While results on the uniform convergence of objective values provide important insights, they do not directly characterize the convergence behaviors of stationary points formulated via subdifferentials. Analyzing the uniform convergence of these set-valued subdifferential mappings of the empirical risk is crucial for a comprehensive understanding of statistical estimation and the behaviors of numerical algorithms in stochastic programming [2, Section 7.2]. Despite its importance, this characterization of subdifferential set convergence remains elusive, warranting further investigation (see Section 1.1 for related works).
@@
This work addresses this gap by introducing a new perspective to understand the uniform convergence behavior of subdifferential sets, specifically for stochastic weakly convex objectives—a prominent class of nonsmooth, nonconvex objectives with widespread applications [3, 4, 5]. Our approach hinges on a fundamental property of weakly convex functions. Specifically, for any pair of weakly convex functions, we establish that over any arbitrary open set within their domains, the uniform bound on the distance between their respective set-valued subdifferential mappings, measured via the Hausdorff metric, is upper bounded by the uniform bound on the distance between any corresponding vector-valued subgradient mappings of these functions (Theorem 1).  This deterministic result immediately implies a crucial connection between the uniform convergence of subdifferentials and the uniform convergence of subgradient mappings in sample average approximations, enabling the analysis of the former through the latter (Theorem 2).
@@
We further apply our technique to an important subclass of stochastic weakly-convex objectives called stochastic convex-composite objectives [6, 7], which include robust phase retrieval [8], blind deconvolution [9], matrix sensing [10, 11], conditional Value-at-Risk [12], and nonlinear quantile regressions [13], commonly encountered in the field of machine learning, statistics, imaging, and risk management (Theorem 3). We establish uniform convergence of subdifferential sets of stochastic weakly-convex objectives by analyzing the convergence behavior of stochastic subgradient mappings (Theorem 4 and Theorem 5). Our result, which does not rely on a common distributional assumption in the literature requiring the subdiffrential mappings to be continuous under Hausdorff metric [1, Theorem 3], provides tight convergence rates in the context of many existing studies (see Section 1.1 for a detailed explanation). For instance, in robust phase retrieval [8, 14], our results (Theorem 6) achieve a tight $\sqrt{d/m}$ rate (modulo logarithmic factors on $d,m$) for uniform convergence of subdifferential mappings under the Hausdorff distance where $d$ is the dimension and $m$ is the sample size. This is complementary to the state-of-the-art $4]{d/m}$ rate, which is established under the graphical distance of subdifferentials, through a different technique based on Attouch’s epigraphical convergence theorem [15, 16]. This leads to new understanding of the nonsmooth landscape of these objectives, e.g., the landscape of the finite sample robust phase retrieval initially studied in [14], including the locations and convergence properties of stationary points.
@@ 1.1 Related Work and Contributions
A significant body of literature has utilized set-valued variational analysis [17] and random set theory [18] to study the uniform convergence of subdifferentials from empirical to population risks [1]. This technique is popular because subdifferential mappings of many nonsmooth and nonconvex functions—including weakly convex functions—often exhibit structural properties such as outer semicontinuity, and the subdifferentials themselves form convex sets, unlike discontinuous subgradient mappings [17]. This technique discretizes the parameter space, achieving initial subdifferential convergence at discrete points via Artstein and Vitale’s strong law of large numbers for random convex sets [19]. The core assumption is then the subdifferential mapping of both population and the finite sample risks are continuous under the Hausdorff metric. This continuity—not merely outer semicontinuity—is crucial, see [1, Remark 2], as it enables the extension of the initial convergence from discrete points to uniform convergence of subdifferential sets over the entire parameter space. Without continuity, using outer-semicontinuity this approach can still ensure uniform convergence of enlarged empirical subdifferentials to enlarged population subdifferentials; however, this does not apply to the original subdifferentials [1].
@@
Therefore, a common and core assumption of this approach to attain uniform convergence is the continuity needed for both the population and empirical subdifferentials; while it facilitates many applications, it is sometimes restrictive and does not capture the full picture of subdifferential convergence [1, Remark 2]. For achieving uniform convergence with a rate guarantee, stronger assumptions, such as Hölder continuity of subdifferentials or related $H$-calmness conditions, are often necessary [20]. Some research aims to relax the requirement to piecewise continuity of subdifferential sets over disjoint regions, but this still necessitates the underlying probability measure being nonatomic to ensure that samples almost surely avoid the discontinuity boundaries of the subdifferentials [21].
@@
This work presents a novel perspective to establish uniform convergence of subdifferential mappings that distinguishes it apart from the existing literature. Specifically, our work targets the domain of stochastic weakly-convex classes, pivotal to numerous data-driven applications [5]. The fundamental principle of our approach is that, for stochastic weakly convex objectives, achieving uniform convergence—and the associated rates—of subdifferential sets using sample average approximation requires only the uniform convergence and rates of any chosen pair of the subgradient mappings of the corresponding empirical and population risk functions. As a result, we can derive uniform convergence (and also rates) for a specific subclass termed stochastic convex-composite objectives, without the need for continuity assumptions on the subdifferential mappings of either population or finite sample objectives often required in the existing approach. Importantly, obtaining uniform convergence for discontinuous subgradient mappings in stochastic convex-composite objectives still demands sophisticated tools from statistical learning theory, such as chaining, the concept of shattering, and the notion of Vapnik–Chervonenkis dimension [22].
@@
Notably, there is a recent advance in the literature that also targets at establishment of uniform subdifferential convergence within the domain of stochastic weakly convex classes [14, 16]. However, there are significant differences between our results and their results. First, there is a subtle difference in terms of the criteria for convergence. Our notion of uniform convergence on any open domain implies their notion of graphical convergence of subdifferential mappings on the same domain, but the reverse is not generally true (see Remark of Theorem 1). Therefore, our results are established under a topologically weaker notion of convergence. Second, there is an important difference in the approaches and the established rates. Their approach, grounded in the utilization of Moreau envelope smoothing [23], and Attouch’s epigraphical convergence theorem [15], achieves a convergence rate whose dependence on the sample size $m$ is often at $m^{-1/4}$. The reason is that the core principle behind their approach, which aims to derive graphical subdifferential distance bounds through the closeness of objective values, often inherently results in suboptimal guarantees in terms of the sample size [16, Section 5]. Our results are notably precise regarding the sample size $m^{-1/2}$, although they apply mainly to stochastic convex-composite objectives and introduce additional logarithmic factors. This indicates that further research is essential to fully delineate the advantages and limitations of both approaches.
@@
Finally, there are many interesting alternative ideas in the literature based on directly smoothing the nonsmooth objective to attain uniform convergence of gradients of the smoothed objectives [24], also used extensively in the statistics literature, e.g., in the context of quantile regressions [25]. These results, however, do not directly characterize the convergence behavior of the original subdifferential sets.
@@ 2 Notation and Basic Definitions
For a function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, its epigraph is the set $\{(x,\alpha)\in\mathbb{R}^{d}\times\mathbb{R}\mid\alpha\geq f(x)\}$. It is called closed if its epigraph is closed in $\mathbb{R}^{d}\times\mathbb{R}$. It is called subdifferentially regular at $x$ if $f(x)$ is finite and its epigraph is Clarke regular at $(x,f(x))$ as a subset of $\mathbb{R}^{d}\times\mathbb{R}$ [17, Definition 7.25]. A function $f:\mathbb{R}^{d}\to\mathbb{R}$ is said to be $\mathcal{C}^{1}$ smooth on $\mathbb{R}^{d}$ if it has continuous gradients $\nabla f:\mathbb{R}^{d}\to\mathbb{R}^{d}$.
@@ 2.1 Weakly Convex Functions and Subdifferentials
We say $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is $\lambda$ locally weakly-convex near $x$ (also known as lower-$\mathcal{C}^{2}$ [17] or semiconvex [26]) if there exists $\epsilon>0$ such that
@@
is convex [17, Chapter 10.G]. We say a function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is a locally weakly convex function on $\mathcal{O}\subseteq\mathbb{R}^{d}$ if at every $x\in\mathcal{O}$, there is $\lambda<\infty$ such that $f$ is $\lambda$ locally weakly-convex near $x$. For a function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ and a point $x$ with $f(x)$ finite, we let $\partial f(x)$ denote the Fréchet subdifferential (or regular [17, Chapter 8.B]) of $f$ at the point $x$,
@@
Here, any $g\in\partial f(x)$ is referred to as a subgradient of $f$ at the point $x$. We say $v$ is a horizon subgradient of $f$ at $x$ with $f(x)$ finite, written as $v\in\partial^{\infty}f(x)$ if there exist sequences $x_{i},v_{i}\in\partial f(x_{i})$ and $\tau_{i}\to 0$ satisfying $(x_{i},f(x_{i}),\tau_{i}v_{i})\to(x,f(x),v)$ [17, Chapter 8.B].
Notably, when $f$ is smooth, the subdifferential $\partial f(x)$ consists only of the gradient $\{\nabla f(x)\}$, while for convex function, it coincides with the subdifferential in convex analysis [27]. For a weakly convex function $f$, the subdifferential $\partial f(x)$ is non-empty, compact, and convex for every $x$ in the interior of ${\rm dom}f$. Notably, these follow from the corresponding results for convex functions [17, Chapter 8].
@@ 2.2 Set-valued Analysis
Our definitions follow the references of Rockafellar and Wet [17], and Aubin and Frankowska [28]. For a set $A\subseteq\mathbb{R}^{d}$, we denote $\left\|{A}\right\|=\sup_{a\in A}\left\|{a}\right\|$.
@@
$\liminf_{n}A_{n}=\{y:\exists y_{n}\in A_{n}~{}\text{s.t.}~{}y_{n}\to y~{}\text{as}~{}n\to\infty\}.$
@@
For every weakly convex function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, the subgradient mapping $\partial f:{\rm int}~{}{\rm dom}(f)\rightrightarrows\mathbb{R}^{d}$ is outersemicontinuous [17, Chapter 8].
@@ 2.3 Probability Space and Random Set-Valued Map
Let $(\Xi,\mathscr{G},\mathbb{P})$ denote a probability space. We frequently consider random set-valued mapping of the form $\mathcal{A}:X\times\Xi\rightrightarrows\mathbb{R}^{d}$. Let $\mathscr{B}$ denote the space of nonempty, compact subsets of $\mathbb{R}^{d}$. By [17, Theorem 14.4], we say the mapping $\xi\mapsto\mathcal{A}(x,\xi)$ is measurable if and only if for every $\mathcal{S}\in\mathscr{B}$, $\mathcal{A}(x,\cdot)^{-1}\mathcal{S}=\{\xi:A(x,\xi)\in\mathcal{S}\}$ is a measurable set under $(\Xi,\mathscr{G},\mathbb{P})$.
By a selection of the random set $\mathcal{A}(x,\xi)$, we refer to a random vector $a(x,\xi)\in\mathcal{A}(x,\xi)$, meaning that $\xi\mapsto a(x,\xi)$ is measurable. Note that such selection exists if $\xi\mapsto\mathcal{A}(x,\xi)$ is closed-valued and measurable [29], see also [28, Theorem 8.4]. For a map $\mathcal{A}:X\times\Xi\rightrightarrows\mathbb{R}^{d}$ and a probability measure $\mathbb{P}$, we define its expectation, following Aumann [29]. At every $x\in X$:
@@
We say $f:\mathbb{R}^{d}\times\Xi\to\mathbb{R}\cup\{+\infty\}$ is a random function when $\xi\mapsto f(x,\xi)$ is measurable at every $x\in X$ [2, Chapter 7]. Suppose a random function $f$ also satisfies $x\mapsto f(x,\xi)$ is locally weakly convex, and real-valued near a point $x\in\mathbb{R}^{d}$, then the subdifferential $\partial f(x,\xi)$ satisfies $\xi\mapsto\partial f(x,\xi)$ is measurable at that $x\in\mathbb{R}^{d}$, following [2, Section 7.2.6] and [24, Proposition 2.1].